OpenStudy (anonymous):

discuss the values ​​of these two equations: 1) x ²-2xsina-cos ² a = 0 2) ² x ²-2x + cos a = 0

OpenStudy (anonymous):

you should play around discriminants of quadratics

OpenStudy (anonymous):

for example first one\[x^2-2x \sin a- \cos^2 a = 0\]\[\Delta=(-2 \sin a)^2-4(-\cos^2 a)=4\sin^2 a +4 \cos^2 a=4>0\]it has two real solutions

OpenStudy (anonymous):

you do the second one :) and tell us what you get

OpenStudy (anonymous):

good night from here :-)

OpenStudy (anonymous):

delta=4-4(cos²x)=4-4cos²x=4(1-cos²x)=4(1-cosx)(1+cosx).